Household Income day A 2016 poll asked people how often they used the Internet. The data collected is included in the table ta) Under the usual null hypothesis, what are the expected values? b) Compute the x2 statistic.
Answer #1

Solution:

Here, we have to use chi square test for independence of two categorical variables.

Null hypothesis: H0: Two categorical variables internet use and household income are independent.

Alternative hypothesis: Ha: Two categorical variables internet use and household income are dependent.

We assume level of significance = α = 0.05

Test statistic formula is given as below:

Chi square = ∑[(O – E)^2/E]

Where, O is observed frequencies and E is expected frequencies.

E = row total * column total / Grand total

We are given

Number of rows = r = 5

Number of columns = c = 3

Degrees of freedom = df = (r – 1)*(c – 1) = 4*2 = 8

α = 0.05

Critical value = 15.50731

(by using Chi square table or excel)

Calculation tables for test statistic are given as below:

Observed Frequencies

How often do you use the Internet?

Household Income

Every day

At least once a week but not every day

Once a week or less

Total

Less than $25,000

105

46

21

172

$25,000 to $49,999

147

35

17

199

$50,000 to $84,999

211

27

11

249

$85,000 to $124,999

213

25

7

245

$125,000 or more

161

17

1

179

Total

837

150

57

1044

Expected Frequencies

How often do you use the Internet?

Household Income

Every day

At least once a week but not every day

Once a week or less

Total

Less than $25,000

137.8966

24.71264

9.390805

172

$25,000 to $49,999

159.5431

28.59195

10.86494

199

$50,000 to $84,999

199.6293

35.77586

13.59483

249

$85,000 to $124,999

196.4224

35.20115

13.37644

245

$125,000 or more

143.5086

25.71839

9.772989

179

Total

837

150

57

1044

Calculations

(O - E)

-32.8966

21.28736

11.6092

-12.5431

6.408046

6.135057

11.37069

-8.77586

-2.59483

16.57759

-10.2011

-6.37644

17.49138

-8.71839

-8.77299

(O - E)^2/E

7.84779

18.33683

14.35164

0.986125

1.436175

3.464255

0.647663

2.15273

0.495271

1.399109

2.956251

3.039595

2.131916

2.955486

7.875311

Test Statistic = Chi square = ∑[(O – E)^2/E] = 70.07614

χ2 statistic = 70.07614

P-value = 0.0000

(By using Chi square table or excel)

P-value < α = 0.05

So, we reject the null hypothesis

There is not sufficient evidence to conclude that two categorical variables internet use and household income are independent.

Part a)

Answer:

The required expected values are given as below:

Expected Frequencies

How often do you use the Internet?

Household Income

Every day

At least once a week but not every day

Once a week or less

Total

Less than $25,000

137.9

24.7

9.4

172

$25,000 to $49,999

159.5

28.6

10.9

199

$50,000 to $84,999

199.6

35.8

13.6

249

$85,000 to $124,999

196.4

35.2

13.4

245

$125,000 or more

143.5

25.7

9.8

179

Total

837

150

57

1044

Part b)

Answer:

χ2 statistic = 70.07614